Fontaine and Monroe are forming a partnership. Fontaine invests a building that has a market value of $336,000; the partnership assumes responsibility for a $118,000 note secured by a mortgage on the property. Monroe invests $93,000 in cash and equipment that has a market value of $68,000. For the partnership, the amounts recorded for Fontaine's Capital account and for Monroe's Capital account are:
Answer #1

For the partnership, the amounts recorded for Fontaine's Capital account and for Monroe's Capital account are:

Fontaine's capital Account = 336000-118000 = $218000

Monroe's capital account = 93000+68000 = $161000